# When Person Re-Identification Meets Event Camera: A Benchmark Dataset and An Attribute-guided Re-Identification Framework

### v1

@@ 1 Introduction
To address the aforementioned issues, some researchers resort to the bio-inspired event cameras for person re-identification [2, 9, 24], due to their advantages in low energy consumption, high dynamic range, no motion blur, and spatial sparsity [12], as shown in Fig. 1. Specifically, Ahmad et al. [2] propose a person re-identification method through event anonymization, aiming to recognize individuals based on behavior and dynamic features without relying on identity information. Cao et al. [9] present an event-guided person re-identification method that leverages sparse-dense complementary learning, combining sparse event data and dense image features to enhance the robustness and accuracy of ReID across different scenarios. Li et al. [24] propose a video person re-identification method that enhances performance by leveraging cross-modality fusion of visual and event data, along with temporal collaboration to capture dynamic motion in challenging scenarios.
@@
Despite these progress, current event-based person re-identification algorithms are still limited by the following issues: 1). Existing person ReID algorithms are typically trained and evaluated on small-scale or simulated event camera datasets, making it difficult to assess their real identification performance and generalization ability. 2). Current ReID algorithms mainly focus on learning event stream features [24, 2] or fusing RGB and event features [9], but fail to consider the semantic information, such as pedestrian attributes long hair, wearing glasses [44], they can only achieve sub-optimal performance. Therefore, it is natural to raise the following question: “How can we design a more effective and generalizable event-based person re-identification framework that leverages not only large-scale, real-world data but also incorporates rich semantic information such as pedestrian attributes?
@@
Considering that the datasets for person re-identification based on event cameras are still scarce and limited in scale, for example, the Event ReID [2] dataset contains only 16,000 samples involving 33 pedestrians. In this paper, we first propose a new benchmark dataset for event stream-based person re-identification to bridge the data gap, termed EvReID. It contains 118,988 image pairs (7 times larger than the current real ReID dataset Event-ReID [2]) and covers 1200 pedestrian identities (36 times more than Event-ReID), with data collected across multiple seasons, scenes, and lighting conditions. We also evaluate 15 state-of-the-art person ReID algorithms, laying a solid foundation for future research in terms of both data and benchmarking. Some representative samples of our EvReID dataset are provided in Fig. 2.
@@
Based on our newly proposed benchmark dataset, we further propose a new pedestrian attribute-guided contrastive learning framework to enhance feature learning for person re-identification, termed TriPro-ReID. This framework not only effectively explores the visual features from both RGB frames and event streams, but also fully utilizes pedestrian attributes as mid-level semantic features. Specifically, given multi-modal input data, we first perform patching and projection to obtain RGB and event tokens, which are then fed into separate ViT backbone networks to learn spatio-temporal features. To facilitate feature interaction between the two modalities, we introduce a cross-modal prompt projector for cross-modal feature fusion. Meanwhile, we employ the VTFPAR++ [48] to predict pedestrian attributes from the input samples and use a text encoder to generate attribute semantic tokens. Subsequently, an attribute prompt injector is introduced to map these semantic tokens and fuse them with the visual features. We adopt widely-used metrics, including the triplet loss, ID loss, and the vision-attribute contrastive loss, to jointly guide the multi-modal feature learning for ReID. An overview of our proposed RGB-Event based person re-identification framework can be found in Fig. 3.
@@ 2 Related Works
In this work, we propose to review the related works on Event-based Vision, Person Re-Identification, and Benchmark Datasets for Video-based Person Re-Identification. More works can be found in the following surveys [12].
@@ 2.1 Event-based Vision
Most computer vision systems rely on RGB modalities, which typically achieve superior performance. However, these systems tend to struggle in low-light environments. To address the shortcomings of RGB-based vision, numerous Event-based methods have been proposed, inspired by the unique pulse-like characteristics of events. In the re-identification field, Ahmad et al. [2] introduce the first Event-based person re-identification (ReID) dataset and develop an Event-anonymous network for privacy-preserving Event-based ReID. In the object detection field, RVT [13], GET [34], and SAST [35] achieve outstanding performance by leveraging a Transformer-based backbone, combined with an LSTM network and specialized attention mechanisms. MvHeat-DET [47] uses a novel heat conduction-based foundational model to fully harness the advantages of events, achieving an optimal balance between accuracy and speed. SpikeYOLO [32], SpikingYOLO [22], EMS-YOLO [39] employ an SNN network to model Event stream data and integrate it with YOLO for object detection. In tracking field, Wang et al. [46], Zhu et al. [60] utilize RGB and Event dual-modal data for tracking and introduce a new multimodal tracking dataset. Mamba-FETrack [19] uses Mamba backbone networks to extract the features to realize more efficient tracking. However, even though Event-based methods address many of the limitations inherent in RGB cameras, RGB cameras still offer richer visual information, which Event cameras lack. This gap often results in Event-based methods achieving lower accuracy compared to their RGB counterparts. Therefore, we propose a dual-branch RGB-Event architecture to leverage the complementary strengths of both modalities, thereby improving performance.
@@ 2.2 Person Re-Identification
Person re-identification (ReID) [38] has received increasing attention, particularly in video-based scenarios where temporal coherence between frames is exploited for more robust matching. Early methods typically adopted CNNs as backbones, such as AP3D [14], which utilizes 3D convolutions to jointly model spatial-temporal cues, and TCLNet [16], which extracts complementary features through temporal modules. CTL [27], PSTA [49], and STRF [5] further enhance temporal modeling by integrating LSTM networks. Graph-based approaches have also been explored, with MGH [53] introducing hypergraph structures to capture high-order spatial-temporal relations, and Chen et al. [10] employing keypoint-based GCNs for refined motion representation. Meanwhile, Transformer-based architectures have shown strong potential for global context modeling. Works like DCCT [30], VDT [56], PFD [43], and SBM [7] adopt transformer backbones to improve representation quality. With the emergence of event cameras, methods such as [4, 3, 1, 9] have explored event stream data to achieve high accuracy under challenging conditions while preserving privacy. Recently, the vision-language model CLIP [36] has demonstrated strong transferability, prompting its adoption in ReID. Li et al. [26] introduce a two-stage training strategy to align image features with textual supervision, while SVLL-ReID [40] generates diverse textual queries to enhance CLIP’s matching capacity. TF-CLIP [54] replaces static text with learned sequence representations, and CLIMB [55] combines CLIP with the Mamba architecture to balance accuracy and efficiency. Several works further incorporate external semantic knowledge. IDEA [51] generates structured captions for cross-modal consistency, and LATex [18] integrates attribute-based prompts to bridge the aerial-ground domain gap. Despite these advances, most existing methods either ignore attribute semantics or rely on heavy models for semantic enhancement. In contrast, we propose an attribute-guided dual-modality ReID framework that leverages lightweight attribute recognition to predict semantic descriptions, encodes them into text features via a text encoder, and fuses them with visual representations. A cross-modal prompting mechanism is further introduced to enable effective interaction and mutual enhancement between modalities, resulting in fine-grained, semantically enriched representations for robust cross-modal ReID.
@@ 2.3 Benchmark Datasets for Video-Based Person Re-Identification
The most widely used datasets for video-based person re-identification (ReID) include PRID-2011 [15], iLIDS-VID [42], MARS [57], DukeMTMC-VideoReID [52], and LS-VID [23]. PRID-2011 comprises 400 video sequences of 200 individuals captured by two static cameras, while iLIDS-VID contains 600 sequences of 300 pedestrians recorded under surveillance conditions from two disjoint camera views. As a large-scale benchmark, MARS is collected from six cameras, offering 17,503 tracklets for 1,261 identities along with 3,248 distractor sequences. DukeMTMC-VideoReID is a subset of the DukeMTMC [37] dataset, consisting of 702 identities for training, 702 for testing, and an additional 408 distractors. LS-VID further scales up video-based ReID with 3,772 annotated identities, offering more diverse visual scenarios. In addition to RGB-based datasets, the event modality has been introduced into the ReID domain. Ahmad et al. [2] proposed the first event-based ReID dataset, Event-ReID, which contains event sequences for 33 individuals. However, existing datasets are either limited to a single modality or constrained by scale, and to date, no real-world person ReID dataset simultaneously includes both RGB and event modalities under a unified setting. To address this gap, we introduce EvReID, a new benchmark designed for dual-modality video-based ReID. It enables comprehensive evaluation of RGB, event, and RGB-Event fusion methods under consistent spatio-temporal conditions, and serves as a foundation for developing and benchmarking future multi-modal person ReID approaches.
@@
01 PRID-2011 [15] 2011 200 40,033 RGB Single Daytime
02 SAIVT-Softbio [8] 2012 152 64,472 RGB Single Daytime
03 iLIDS-VID [42] 2014 300 42,460 RGB Single Daytime
04 HDA Person [33] 2014 53 2,976 RGB Single Daytime
05 MARS [57] 2016 1,261 1,067,516 RGB Single Daytime
06 DukeMTMC-VideoReID [52] 2018 1,404 815,420 RGB Single Daytime
07 LS-VID [23] 2019 3,772 2,982,685 RGB Single Day, Night
08 Event ReID [2] 2023 33 16,000 Event Single Daytime
@@ 3.2 Data Collection and Statistical Analysis
As shown in Table 1, we compare several representative video-based person re-identification (ReID) datasets with our proposed EvReID dataset. Most existing benchmarks are based solely on the RGB modality, which performs well under well-lit conditions but struggles to capture rapid motion or subtle visual cues in low-light or nighttime environments. To address these limitations, recent efforts have attempted to simulate event-based reID by generating pseudo-event data from RGB frames. Notably, Ahmad et al. [2] proposed the first real-world Event-based ReID dataset, emphasizing privacy protection and robustness in complex scenes. However, their dataset contains only 33 identities, which significantly limits its scalability and suitability for large-scale ReID research. Moreover, many existing datasets are collected in overly constrained or homogeneous environments, limiting their generalization to real-world deployment scenarios.
@@ 3.3 Benchmark Baselines
To establish a comprehensive benchmark dataset for RGB-Event based person re-identification, we select 15 SOTA or representative methods for evaluation on our proposed dataset, including: 1). CNN-based: OSNet [58], TCLNet [16], AP3D [14], MGH [53], STMN [11], PSTA [49], GRL [29], BiCnet-TKS [17], SINet [6], and SDCL [9]. 2). Transformer-based: DCCT [30], CLIP-ReID [25], TF-CLIP [54], DeMo [50] and CLIMB-ReID [55].
@@ 4.1 Preliminary: CLIP and CLIP-ReID
CLIP [36] consists of a visual encoder $\mathcal{V(\cdot)}$ and a text encoder $\mathcal{T(\cdot)}$, which are jointly trained using contrastive learning to project input images and texts into a shared representation space. Specifically, let ${\{img_{i},text_{i}\}_{i=1}^{B}}$ denote a batch of $B$ paired visual-language training samples, where $\{img_{i}\}$ is an image and $\{text_{i}\}$ is the corresponding textual description. CLIP employs the two encoders along with two linear projection layers to encode images and texts into corresponding feature embeddings. The similarity between image and text features is computed as:
@@
where $\mathcal{J}_{v}$ and $\mathcal{J}_{t}$ are linear layers that project features into a unified embedding space. Two contrastive losses, denoted as $L{v2t}$ and $L{t2v}$, are employed to align visual and textual features during training. In downstream tasks, the textual input is typically formulated as prompts such as “A photo/video of a {class},” where “{}” is replaced with a specific class label (e.g., cat). However, in person ReID, identity annotations are merely represented as integer indices rather than meaningful text. To bridge this gap, CLIP-ReID [26] introduces learnable identity-specific tokens to construct personalized textual prompts. These prompts take the form of “A photo of a $[X]_{1},[X]_{2},\cdots,[X]_{Num}$ person”. During an additional training stage, the parameters of $\mathcal{V}(\cdot)$ and $\mathcal{T}(\cdot)$ are frozen, and the identity tokens are optimized using $L{v2t}$ and $L_{t2v}$ to learn a discriminative but implicit textual representation for each identity. While CLIP-ReID achieves promising results on conventional RGB-based person ReID datasets, a simple dual-branch CLIP-ReID architecture proves insufficient when applied to RGB-Event dual-modality scenarios. It fails to fully leverage the complementary temporal and modality-specific cues provided by the Event stream. More critically, text prompts such as “A photo of a $[X]_{1},[X]_{2},\cdots,[X]_{Num}$ person”, though identity-aware, do not explicitly model or utilize the fine-grained pedestrian attribute information inherently present in video frames.

### v2

@@ 1 Introduction
To address the aforementioned issues, some researchers resort to the bio-inspired event cameras for person re-identification [Ahmad_2023_ICCV, cao2023event, li2025event], due to their advantages in low energy consumption, high dynamic range, no motion blur, and spatial sparsity [gallego2020eventSurvey], as shown in Fig. 1. Specifically, Ahmad et al. [Ahmad_2023_ICCV] propose a person re-identification method through event anonymization, aiming to recognize individuals based on behavior and dynamic features without relying on identity information. Cao et al. [cao2023event] present an event-guided person re-identification method that leverages sparse-dense complementary learning, combining sparse event data and dense image features to enhance the robustness and accuracy of ReID across different scenarios. Li et al. [li2025event] propose a video person re-identification method that enhances performance by leveraging cross-modality fusion of visual and event data, along with temporal collaboration to capture dynamic motion in challenging scenarios.
@@
Despite these progress, current event-based person re-identification algorithms are still limited by the following issues: 1). Existing person ReID algorithms are typically trained and evaluated on small-scale or simulated event camera datasets, making it difficult to assess their real identification performance and generalization ability. 2). Current ReID algorithms mainly focus on learning event stream features [li2025event, Ahmad_2023_ICCV] or fusing RGB and event features [cao2023event], but fail to consider the semantic information, such as pedestrian attributes long hair, wearing glasses [wang2022PARSurvey], they can only achieve sub-optimal performance. Therefore, it is natural to raise the following question: “How can we design a more effective and generalizable event-based person re-identification framework that leverages not only large-scale, real-world data but also incorporates rich semantic information such as pedestrian attributes?
@@
Considering that the datasets for person re-identification based on event cameras are still scarce and limited in scale, for example, the Event ReID [Ahmad_2023_ICCV] dataset contains only 16,000 samples involving 33 pedestrians. In this paper, we first propose a new benchmark dataset for event stream-based person re-identification to bridge the data gap, termed EvReID. It contains 118,988 image pairs (7 times larger than the current real ReID dataset Event-ReID [Ahmad_2023_ICCV]) and covers 1200 pedestrian identities (36 times more than Event-ReID), with data collected across multiple seasons, scenes, and lighting conditions. We also evaluate 15 state-of-the-art person ReID algorithms, laying a solid foundation for future research in terms of both data and benchmarking. Some representative samples of our EvReID dataset are provided in Fig. 2.
@@
Based on our newly proposed benchmark dataset, we further propose a new pedestrian attribute-guided contrastive learning framework to enhance feature learning for person re-identification, termed TriPro-ReID. This framework not only effectively explores the visual features from both RGB frames and event streams, but also fully utilizes pedestrian attributes as mid-level semantic features. Specifically, given multi-modal input data, we first perform patching and projection to obtain RGB and event tokens, which are then fed into separate ViT backbone networks to learn spatio-temporal features. To facilitate feature interaction between the two modalities, we introduce a cross-modal prompt projector for cross-modal feature fusion. Meanwhile, we employ the VTFPAR++ [wang2024spatio] to predict pedestrian attributes from the input samples and use a text encoder to generate attribute semantic tokens. Subsequently, an attribute prompt injector is introduced to map these semantic tokens and fuse them with the visual features. We adopt widely-used metrics, including the triplet loss, ID loss, and the vision-attribute contrastive loss, to jointly guide the multi-modal feature learning for ReID. An overview of our proposed RGB-Event based person re-identification framework can be found in Fig. 3.
@@ 2 Related Works
In this work, we propose to review the related works on Event-based Vision, Person Re-Identification, and Benchmark Datasets for Video-based Person Re-Identification. More works can be found in the following surveys [gallego2020eventSurvey].
@@ 2.1 Event-based Vision
Most computer vision systems rely on RGB modalities, which typically achieve superior performance. However, these systems tend to struggle in low-light environments. To address the shortcomings of RGB-based vision, numerous Event-based methods have been proposed, inspired by the unique pulse-like characteristics of events. In the re-identification field, Ahmad et al. [Ahmad_2023_ICCV] introduce the first Event-based person re-identification (ReID) dataset and develop an Event-anonymous network for privacy-preserving Event-based ReID. In the object detection field, RVT [gehrig2023recurrent], GET [peng2023get], and SAST [peng2024scene] achieve outstanding performance by leveraging a Transformer-based backbone, combined with an LSTM network and specialized attention mechanisms. MvHeat-DET [wang2024EvDET200K] uses a novel heat conduction-based foundational model to fully harness the advantages of events, achieving an optimal balance between accuracy and speed. SpikeYOLO [luo2024spikeYOLO], SpikingYOLO [kim2020spikingyolo], EMS-YOLO [su2023emsyolo] employ an SNN network to model Event stream data and integrate it with YOLO for object detection. In tracking field, Wang et al. [wang2024long], Zhu et al. [zhu2024crsot] utilize RGB and Event dual-modal data for tracking and introduce a new multimodal tracking dataset. Mamba-FETrack [huang2024mamba] uses Mamba backbone networks to extract the features to realize more efficient tracking. However, even though Event-based methods address many of the limitations inherent in RGB cameras, RGB cameras still offer richer visual information, which Event cameras lack. This gap often results in Event-based methods achieving lower accuracy compared to their RGB counterparts. Therefore, we propose a dual-branch RGB-Event architecture to leverage the complementary strengths of both modalities, thereby improving performance.
@@ 2.2 Person Re-Identification
Person re-identification (ReID) [shu2021large] has received increasing attention, particularly in video-based scenarios where temporal coherence between frames is exploited for more robust matching. Early methods typically adopted CNNs as backbones, such as AP3D [gu2020AP3D], which utilizes 3D convolutions to jointly model spatial-temporal cues, and TCLNet [TCLNet], which extracts complementary features through temporal modules. CTL [liu2021ctl], PSTA [PSTA], and STRF [aich2021strf] further enhance temporal modeling by integrating LSTM networks. Graph-based approaches have also been explored, with MGH [DBLP:conf/cvpr/YanQC0ZT020] introducing hypergraph structures to capture high-order spatial-temporal relations, and Chen et al. [chen2022keypoint] employing keypoint-based GCNs for refined motion representation. Meanwhile, Transformer-based architectures have shown strong potential for global context modeling. Works like DCCT [liu2023deeply], VDT [zhang2024view], PFD [wang2022pose], and SBM [bai2022salient] adopt transformer backbones to improve representation quality. With the emergence of event cameras, methods such as [ahmad2024event, ahmad2023person, ahmad2022event, cao2023event] have explored event stream data to achieve high accuracy under challenging conditions while preserving privacy. Recently, the vision-language model CLIP [radford2021learning] has demonstrated strong transferability, prompting its adoption in ReID. Li et al. [li2023clip] introduce a two-stage training strategy to align image features with textual supervision, while SVLL-ReID [wang2025image] generates diverse textual queries to enhance CLIP’s matching capacity. TF-CLIP [tfclip] replaces static text with learned sequence representations, and CLIMB [yu2025climb] combines CLIP with the Mamba architecture to balance accuracy and efficiency. Several works further incorporate external semantic knowledge. IDEA [wang2025idea] generates structured captions for cross-modal consistency, and LATex [hu2025latex] integrates attribute-based prompts to bridge the aerial-ground domain gap. Despite these advances, most existing methods either ignore attribute semantics or rely on heavy models for semantic enhancement. In contrast, we propose an attribute-guided dual-modality ReID framework that leverages lightweight attribute recognition to predict semantic descriptions, encodes them into text features via a text encoder, and fuses them with visual representations. A cross-modal prompting mechanism is further introduced to enable effective interaction and mutual enhancement between modalities, resulting in fine-grained, semantically enriched representations for robust cross-modal ReID.
@@ 2.3 Benchmark Datasets for Video-Based Person Re-Identification
The most widely used datasets for video-based person re-identification (ReID) include PRID-2011 [hirzer2011person], iLIDS-VID [wang2014person], MARS [zheng2016mars], DukeMTMC-VideoReID [wu2018exploit], and LS-VID [li2019global]. PRID-2011 comprises 400 video sequences of 200 individuals captured by two static cameras, while iLIDS-VID contains 600 sequences of 300 pedestrians recorded under surveillance conditions from two disjoint camera views. As a large-scale benchmark, MARS is collected from six cameras, offering 17,503 tracklets for 1,261 identities along with 3,248 distractor sequences. DukeMTMC-VideoReID is a subset of the DukeMTMC [ristani2016performance] dataset, consisting of 702 identities for training, 702 for testing, and an additional 408 distractors. LS-VID further scales up video-based ReID with 3,772 annotated identities, offering more diverse visual scenarios. In addition to RGB-based datasets, the event modality has been introduced into the ReID domain. Ahmad et al. [Ahmad_2023_ICCV] proposed the first event-based ReID dataset, Event-ReID, which contains event sequences for 33 individuals. However, existing datasets are either limited to a single modality or constrained by scale, and to date, no real-world person ReID dataset simultaneously includes both RGB and event modalities under a unified setting. To address this gap, we introduce EvReID, a new benchmark designed for dual-modality video-based ReID. It enables comprehensive evaluation of RGB, event, and RGB-Event fusion methods under consistent spatio-temporal conditions, and serves as a foundation for developing and benchmarking future multi-modal person ReID approaches.
@@
01 PRID-2011 [hirzer2011person] 2011 200 40,033 RGB Single Daytime
02 SAIVT-Softbio [bialkowski2012database] 2012 152 64,472 RGB Single Daytime
03 iLIDS-VID [wang2014person] 2014 300 42,460 RGB Single Daytime
04 HDA Person [nambiar2014multi] 2014 53 2,976 RGB Single Daytime
05 MARS [zheng2016mars] 2016 1,261 1,067,516 RGB Single Daytime
06 DukeMTMC-VideoReID [wu2018exploit] 2018 1,404 815,420 RGB Single Daytime
07 LS-VID [li2019global] 2019 3,772 2,982,685 RGB Single Day, Night
08 Event ReID [Ahmad_2023_ICCV] 2023 33 16,000 Event Single Daytime
@@ 3.2 Data Collection and Statistical Analysis
As shown in Table 1, we compare several representative video-based person re-identification (ReID) datasets with our proposed EvReID dataset. Most existing benchmarks are based solely on the RGB modality, which performs well under well-lit conditions but struggles to capture rapid motion or subtle visual cues in low-light or nighttime environments. To address these limitations, recent efforts have attempted to simulate event-based reID by generating pseudo-event data from RGB frames. Notably, Ahmad et al. [Ahmad_2023_ICCV] proposed the first real-world Event-based ReID dataset, emphasizing privacy protection and robustness in complex scenes. However, their dataset contains only 33 identities, which significantly limits its scalability and suitability for large-scale ReID research. Moreover, many existing datasets are collected in overly constrained or homogeneous environments, limiting their generalization to real-world deployment scenarios.
@@ 3.3 Benchmark Baselines
To establish a comprehensive benchmark dataset for RGB-Event based person re-identification, we select 15 SOTA or representative methods for evaluation on our proposed dataset, including: 1). CNN-based: OSNet [zhou2019omniscalefeaturelearningperson], TCLNet [TCLNet], AP3D [gu2020AP3D], MGH [DBLP:conf/cvpr/YanQC0ZT020], STMN [eom2021video], PSTA [PSTA], GRL [liu2021watching], BiCnet-TKS [BiCnet-TKS], SINet [bai2022SINet], and SDCL [cao2023event]. 2). Transformer-based: DCCT [liu2023deeply], CLIP-ReID [li2022clip], TF-CLIP [tfclip], DeMo [wang2025decoupled] and CLIMB-ReID [yu2025climb].
@@ 4.1 Preliminary: CLIP and CLIP-ReID
CLIP [radford2021learning] consists of a visual encoder $\mathcal{V(\cdot)}$ and a text encoder $\mathcal{T(\cdot)}$, which are jointly trained using contrastive learning to project input images and texts into a shared representation space. Specifically, let $\{img_{i},text_{i}\}_{i=1}^{B}$} denote a batch of $B$ paired visual-language training samples, where $\{img_{i}\}$ is an image and $\{text_{i}\}$ is the corresponding textual description. CLIP employs the two encoders along with two linear projection layers to encode images and texts into corresponding feature embeddings. The similarity between image and text features is computed as:
@@
where $\mathcal{J}_{v}$ and $\mathcal{J}_{t}$ are linear layers that project features into a unified embedding space. Two contrastive losses, denoted as Lv2tL{v2t} and Lt2vL{t2v}, are employed to align visual and textual features during training. In downstream tasks, the textual input is typically formulated as prompts such as “A photo/video of a {class},” where “{}” is replaced with a specific class label (e.g., cat). However, in person ReID, identity annotations are merely represented as integer indices rather than meaningful text. To bridge this gap, CLIP-ReID [li2023clip] introduces learnable identity-specific tokens to construct personalized textual prompts. These prompts take the form of “A photo of a $X]_{1},[X]_{2},\cdots,[X]_{Num}$ person”. During an additional training stage, the parameters of $\mathcal{V}(\cdot)$ and $\mathcal{T}(\cdot)$ are frozen, and the identity tokens are optimized using Lv2tL{v2t} and $L_{t2v}$ to learn a discriminative but implicit textual representation for each identity. While CLIP-ReID achieves promising results on conventional RGB-based person ReID datasets, a simple dual-branch CLIP-ReID architecture proves insufficient when applied to RGB-Event dual-modality scenarios. It fails to fully leverage the complementary temporal and modality-specific cues provided by the Event stream. More critically, text prompts such as “A photo of a $X]_{1},[X]_{2},\cdots,[X]_{Num}$ person”, though identity-aware, do not explicitly model or utilize the fine-grained pedestrian attribute information inherently present in video frames.